# Stochastic Momentum Methods for Non-smooth Non-Convex Finite-Sum Coupled Compositional Optimization

### v1

@@ 3 Preliminaries
[…]
@@
It is known that, for $\lambda\in(0,\rho^{-1})$, $f_{\lambda}(g)$ is $L_{f}:=\frac{2-\lambda\rho}{\lambda-\lambda^{2}\rho}$-smooth [36] and
@@ 4 Smoothing of Non-smooth FCCO
We can easily show that when $g_{i}$ is $L_{g}$-smooth, then $F_{\lambda}(\mathbf{w})$ is $L_{F}$-smooth with $L_{F}=\frac{C_{g}^{2}(2-\lambda\rho_{f})}{\lambda-\lambda^{2}\rho_{f}}+C_{f}L_{g}$ (cf. Lemma A.1). As a result, the problem becomes a smooth FCCO where both inner functions and outer functions are smooth, which makes it possible to employ existing techniques to develop stochastic momentum methods to find an $\epsilon$-stationary point of (3), which is a point $\mathbf{w}$ satisfying $\mathbb{E}[\|\nabla F_{\lambda}(\mathbf{w})\|]\leq\epsilon$. However, what this implies for solving the original problem and what constitutes a good choice of $\lambda$ remain unclear. Below, we present a theory to address these questions. We introduce the following notion of stationarity of the original problem.
@@
Remark: In Assumption 4.3, once $\epsilon$ is smaller than $c$, the lower bound $\delta$ does not depend on $\epsilon$.
@@
Remark: The proof is given in Appendix A.4. It is notable that the Assumption 4.3 is easily verifiable (cf. Appendix F.4 for empirical evidence). In addition, when we consider the applications in non-convex constrained optimization, this condition is commonly used in existing analyses.
@@ 4.1 Single-loop Methods for Smooth Inner Functions
There exist constants $\sigma_{0}\geq 0$ and $\sigma_{1}\geq 0$ such that the following statements hold for $g_{i}(\mathbf{w})$ and $g_{i}(\mathbf{w},\xi_{i})$ for $i=1,\dots,n$ for any $\mathbf{w}\in\mathbb{R}^{d}$ $\mathbb{E}\|g_{i}(\mathbf{w},\xi_{i})-g_{i}(\mathbf{w})\|^{2}\leq\sigma_{0}^{2}$, $\mathbb{E}\left\|\partial g_{i}(\mathbf{w},\xi_{i})-\partial g_{i}(\mathbf{w})\right\|^{2}\leq\sigma_{1}^{2}$.
@@
There exist a constant $C_{g}$ such that $\mathbb{E}_{\xi}\|g_{i}(\mathbf{w},\xi)-g_{i}(\mathbf{w}^{\prime},\xi)\|_{2}]\leq C_{g}\|\mathbf{w}-\mathbf{w}^{\prime}\|_{2}$.
It is notable that these assumptions have been made in [6]. We refer to Assumption 4.6 as mean Lipchitz continuity of zero-order (MLC0). Let $B_{1}$ and $B_{2}$ denote outer and inner batch sizes.
@@
Under Assumption 3.2 (A1), 4.5 and 4.6, by setting $\lambda=\Theta(\epsilon),\beta=\Theta(\min\{B_{1},B_{2}\}\epsilon^{2}),\gamma=\Theta(B_{2}\epsilon^{4}),\eta=\Theta(\frac{B_{1}\sqrt{B_{2}}}{n}\epsilon^{3})$, SONEX with $\gamma^{\prime}=1-\gamma+\frac{n-B_{1}}{B_{1}(1-\gamma)}$ converges to an approximate $\epsilon$-stationary solution of (1) within $T=O(\frac{n}{B_{1}\sqrt{B_{2}}}\epsilon^{-5})$ iterations.
@@ 4.2 A Double-loop Algorithm for Smooth or Weakly-Convex Inner Functions
Recall that $\nabla F_{\lambda,\nu}(\mathbf{w})=\frac{1}{\nu}(\mathbf{w}-\text{prox}_{\nu F_{\lambda}}(\mathbf{w}))$. Given an approximate solution of $\text{prox}_{\nu F_{\lambda}}(\mathbf{w})$, denoted by $\hat{\mathbf{z}}$, of the inner minimization problem in (5), we can use $\frac{1}{\nu}(\mathbf{w}-\hat{\mathbf{z}})$ as an inexact gradient $F_{\lambda,\nu}(\mathbf{w})$ to update $\mathbf{w}$ with a momentum method in order to solve (5). To obtain an estimate $\hat{\mathbf{z}}_{t}$ at the $t$-th iteration, we employ a recent algorithm ALEXR [5] for solving the inner minimax problem. With an estimate $\hat{\mathbf{z}}_{t}$, we will update $\mathbf{w}_{t+1}$ by a momentum update. We present the detailed steps of our double-loop algorithm named ALEXR2 in Algorithm 2, where $U_{\bar{f}_{i}^{*}}(\mathbf{y}_{i},\mathbf{y}_{i}^{\prime})$ is the Bregman divergence induced by $\bar{f}_{i}^{*}$, namely, $U_{\bar{f}_{i}^{*}}(\mathbf{y}_{i},\mathbf{y}_{i}^{\prime}):=\bar{f}_{i}^{*}(\mathbf{y}_{i})-\bar{f}_{i}^{*}(\mathbf{y}_{i}^{\prime})-\left\langle\partial\bar{f}_{i}^{*}(\mathbf{y}_{i}^{\prime}),\mathbf{y}_{i}-\mathbf{y}_{i}^{\prime}\right\rangle$. For the sake of convergence analyses we need $U_{\bar{f}_{i}^{*}}$ to be bounded, as stated in the assumption below. Besides, we would like to point out that the sequences of $y_{i}$ can be updated similar to $u_{i}$ sequences in SONEX similar to [5], i.e., $y_{i,t}=\nabla f_{i,\lambda}(u_{i,t})$, where $u_{i,t+1}=(1-\hat{\gamma})u_{i,t}+\hat{\gamma}g_{i}(\mathbf{z}_{t,k};\mathcal{B}_{i,2}^{t,k})+\hat{\gamma}\theta(g_{i}(\mathbf{z}_{t,k};\mathcal{B}_{i,2}^{t,k})-g_{i}(\mathbf{z}_{t-1,k};\mathcal{B}_{i,2}^{t,k}))$, where $\hat{\gamma}=\gamma/(1+\gamma)$.
@@
Remark: Though looking strong, it is satisfied by many practical convex lipschitz-continuous functions such as hinge function, smoothed hinge function(i.e. $f_{i,\lambda}$ or $\bar{f}_{i}$ in this paper), etc.
@@
5:        Sample $\mathcal{B}_{1}^{t,k}\subset\{1,\dots,n\}$ and two independent batches of data $\mathcal{B}_{i,2}^{t,k}$ and ~$\tilde{\mathcal{B}}_{i,2}^{t,k}$ from the distribution of $\xi_{i}$ for $i\in\mathcal{B}_{1}^{t,k}$.
@@
[…]
@@ 5 Smoothed Hinge Penalty Method for Constrained Optimization
$\min_{\mathbf{w}\in\mathbb{R}^{d}}\Phi_{\lambda}(\mathbf{w}):=g_{0}(\mathbf{w})+\frac{1}{m}\sum_{i=1}^{m}f_{\lambda}(g_{i}(\mathbf{w}))=g_{0}(\mathbf{w})+\frac{1}{m}\sum_{i=1}^{m}\max_{y_{i}\in[0,\rho]}\left\{y_{i}g_{i}(\mathbf{w})-\frac{\lambda}{2}y_{i}^{2}\right\},$ (9)
@@
Suppose Assumption 5.1 hold. If $\rho>\frac{m(C_{g}+1)}{\delta}$ and $\lambda=\frac{\epsilon}{\rho}$, a nearly $\epsilon$-stationary solution $\mathbf{w}$ to (9) is also a nearly $O(\epsilon)$-KKT solution to the original problem (7) in the sense that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$, $\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$, $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon)$.
@@
  • if $\{g_{i}\}_{i=0}^{m}$ are weakly convex, then with $\tilde{\mathcal{O}}\left(\frac{m}{B_{1}B_{2}\delta^{4}\epsilon^{5}}\right)$ iterations ALEXR2 find a nearly $\epsilon$-KKT solution $\mathbf{w}$ for (7) satisfying that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\mathbb{E}\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\mathbb{E}\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$ and it holds with probability [^1] If $|g_{i}(\mathbf{w})|<+\infty$ for any $i$ and $\mathbf{w}$, this high probability result can be replaced by $\mathbb{E}[\max\limits_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})]\leq O(\epsilon)$ and $\mathbb{E}[\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|]\leq O(\epsilon)$. $1-O(\epsilon)$ that $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon)$.
@@
Notably, our method not only improve the rate but also enjoy an improved dependence on $\delta$, compared to $O(\epsilon^{-6}\delta^{-6})$ in prior work [8]. We conduct experiments to verify the effectiveness of the proposed algorithms. For non-smooth non-convex FCCO, we consider GDRO with CVaR divergence [5]. For non-convex constrained optimization, we follow [8] and consider two tasks, namely AUC maximization with ROC fairness constraints and continual learning with non-forgetting constraints.
@@
We use 3 datasets: Camelyon17, Amazon [37], and CelebA [38]. The first two datasets are from WILDS Benchmark for evaluating methods tackling the distributional shift [39], where Camelyon17 has 30 groups and Amazon has 1252 groups. CelebA is a large-scale facial attribute dataset containing over 200,000 celebrity images annotated with 40 binary attributes. We select 4 binary attributes ‘Attractive’, ‘Mouth_Slightly_Open’, ‘Male’ and ‘Blonde_Hair’ and construct 16 groups, where ‘Blonde_Hair’ also serves as the label for classification. We compare SONEX with SONX and OOA, where OOA is a stochastic primal-dual algorithm  [^2] We extend the original OOA [40] to GDRO with CVaR setting as in [5]. We use pretrained Densenet121 [41], Distilbert [42] and ResNet50 [43] for Camelyon17, Amazon and CelebA, respectively. We perform Adam-type update for SONEX. We run each experiment for 3 random seeds and report their average performance. The hyperparameter tuning is discussed in Appendix F.1.
@@
Figure 3: Training curves of Target $\Delta$ACC values (fig 1,3) and constraint violation (fig 2,4) of different methods. The format of label is "Algorithm(penalty function, $\rho$)", and SH, smH mean square hinge and smoothed hinge, respectively.

### v2

@@ 3 Preliminaries
Let $\|\cdot\|$ be the Euclidean norm. For $g(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{1}}$, let $\nabla g(\cdot)\in\mathbb{R}^{d\times d_{1}}$ be its Jacobian and let $\|\nabla g(\cdot)\|=\max_{\mathbf{u}\in\mathbb{R}^{d},\|\mathbf{u}\|=1}\|\nabla g(\cdot)\mathbf{u}\|$ be the operator norm. Let $g_{i}(\mathbf{w},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi\in\mathcal{B}}g_{i}(\mathbf{w},\xi),i=1,2,\cdots,n$ be an estimator of $g_{i}$ based on samples from a minibatch $\mathcal{B}$. A function $f(\cdot)$ is monotonically non-decreasing if for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{d_{1}}$ satisfying that $\mathbf{x}_{1,k}\leq\mathbf{x}_{2,k},k\in[d_{1}]$ where $\mathbf{x}_{k}$ denote the $k$-th element of $\mathbf{x}$, it holds that $f(\mathbf{x}_{1})\leq f(\mathbf{x}_{2})$. A function $f(\cdot)$ is $\rho$-weakly convex for $\rho\geq 0$ if $f(\cdot)+\frac{\rho}{2}\left\|\cdot\right\|^{2}$ is convex. A function $g(\cdot)$ is $L$-smooth for $L\geq 0$ if $g$ is differentiable and $\nabla g(\cdot)$ is $L$-Lipschitz continuous. A vector-valued mapping $g(\cdot)$ is $L$-smooth for $L\geq 0$ if $g$ is differentiable at each component and $\nabla g(\cdot)$ is $L$-Lipschitz continuous w.r.t to the operator norm. Given a $\rho$-weakly convex function $f(\cdot)$, we denote its regular subdifferential as $\partial f(\cdot)$. For a $\rho$-weakly convex function $f(\cdot)$ and a constant $\lambda\in(0,\rho^{-1})$, the proximal mapping and the Moreau envelope of $f(\cdot)$ with parameter $\lambda$ are defined as follows, respectively,
@@
It is known that, for $\lambda\in(0,\rho^{-1})$, $f_{\lambda}(\cdot)$ is $L_{f}:=\max\{\frac{1}{\lambda},\frac{\rho}{1-\lambda\rho}\}$-smooth [33] and
@@
Remark: The Lipschitz continuity of $f$ and $g$ in Assumption 3.2 are fairly standard for FCCO problems. Since we target at non-smooth $f$ and $g$, Lipchitz continuity is a minimal assumption. In the considered applications of GDRO and learning with fairness constraints in section 6, $f$ is hinge and hence Lipschitz continuous.
@@ 4 Smoothing of Non-smooth FCCO
We can easily show that when $g_{i}$ is $L_{g}$-smooth, then $F_{\lambda}(\mathbf{w})$ is $L_{F}$-smooth with $L_{F}=C_{g}^{2}\max\{\frac{1}{\lambda},\frac{\rho_{f}}{1-\lambda\rho_{f}}\}+C_{f}L_{g}$ (cf. Lemma A.1). As a result, the problem becomes a smooth FCCO where both inner functions and outer functions are smooth, which makes it possible to employ existing techniques to develop stochastic momentum methods to find an $\epsilon$-stationary point of (3), which is a point $\mathbf{w}$ satisfying $\mathbb{E}[\|\nabla F_{\lambda}(\mathbf{w})\|]\leq\epsilon$. However, what this implies for solving the original problem and what constitutes a good choice of $\lambda$ remain unclear. Below, we present a theory to address these questions. We introduce the following notion of stationarity of the original problem.
@@
Remark: In Assumption 4.3, once $\epsilon$ is smaller than $c$, the lower bound $\delta$ does not depend on $\epsilon$. The empirical justification for this assumption is provided in subsection F.5.
@@
Remark: The proof is given in Appendix A.4. It is notable that the Assumption 4.3 is easily verifiable (cf. Appendix F.5 for empirical evidence). In addition, when we consider the applications in non-convex constrained optimization, this condition is commonly used in existing analyses [7, 8].
@@ 4.1 Single-loop Methods for Smooth Inner Functions
There exist constants $\sigma_{0}\geq 0$ and $\sigma_{1}\geq 0$ such that the following statements hold for $g_{i}(\mathbf{w})$ and $g_{i}(\mathbf{w},\xi_{i})$: for $i=1,\dots,n$ and any $\mathbf{w}\in\mathbb{R}^{d}$
$\mathbb{E}\|g_{i}(\mathbf{w},\xi_{i})-g_{i}(\mathbf{w})\|^{2}\leq\sigma_{0}^{2},\mathbb{E}\left\|\partial g_{i}(\mathbf{w},\xi_{i})-\partial g_{i}(\mathbf{w})\right\|^{2}\leq\sigma_{1}^{2}$
@@
There exist a constant $C_{g}$ such that $\mathbb{E}_{\xi}\|g_{i}(\mathbf{w},\xi)-g_{i}(\mathbf{w}^{\prime},\xi)\|_{2}\leq C_{g}\|\mathbf{w}-\mathbf{w}^{\prime}\|_{2}$.
It is notable that these assumptions have been made in [6] which is important for analyzing variance reduction technique such as MSVR. We refer to Assumption 4.6 as mean Lipchitz continuity of zero-order (MLC0). Let $B_{1}$ and $B_{2}$ denote outer and inner batch sizes.
@@
Under Assumption 3.2 (A1), 4.5 and 4.6, by setting $\lambda=\Theta(\epsilon),\beta=\Theta(\min\{B_{1},B_{2}\}\epsilon^{2}),\gamma=\Theta(B_{2}\epsilon^{4}),\eta=\Theta(\frac{B_{1}\sqrt{B_{2}}}{n}\epsilon^{3})$, SONEX with $\gamma^{\prime}=1-\gamma+\frac{n-B_{1}}{B_{1}(1-\gamma)}$ and $\gamma\leq\frac{1}{2}$ converges to an approximate $\epsilon$-stationary solution of (1) within $T=O(\frac{n}{B_{1}\sqrt{B_{2}}}\epsilon^{-5})$ iterations.
@@ 4.2 A Double-loop Algorithm for Smooth or Weakly-Convex Inner Functions
[…]
@@
Remark: We point out that this condition is not strong under our setting where $\bar{f}_{i}$ is lipschitz continuous and $dom(\bar{f}_{i}^{*})$ is consequently bounded and it is satisfied by many practical convex lipschitz-continuous functions such as hinge function, smoothed hinge function(i.e. $f_{i,\lambda}$ or $\bar{f}_{i}$ in this paper), etc.
@@
5:   Sample $\mathcal{B}_{1}^{t,k}\subset\{1,\dots,n\}$ and two independent batches of data $\mathcal{B}_{i,2}^{t,k}$ and $\tilde{\mathcal{B}}_{i,2}^{t,k}$ from the distribution of $\xi_{i}$ for $i\in\mathcal{B}_{1}^{t,k}$.
@@
15:  $\text{Update }\mathbf{w}_{t+1}=\mathbf{w}_{t}-\alpha\mathbf{v}_{t+1}\text{ or use Adam-type update}$
@@ 5 Smoothed Hinge Penalty Method for Constrained Optimization
$\hskip-10.84006pt\min_{\mathbf{w}\in\mathbb{R}^{d}}\Phi_{\lambda}(\mathbf{w}):=g_{0}(\mathbf{w})+\frac{1}{m}\sum_{i=1}^{m}f_{\lambda}(g_{i}(\mathbf{w}))=g_{0}(\mathbf{w})+\frac{1}{m}\sum_{i=1}^{m}\max_{y_{i}\in[0,\rho]}\left\{y_{i}g_{i}(\mathbf{w})-\frac{\lambda}{2}y_{i}^{2}\right\},$ (9)
@@
Suppose Assumption 5.1 hold. If $\rho>\frac{m(C_{g}+1)}{\delta}$ and $\lambda=\frac{\epsilon}{\rho}$, a nearly $\epsilon$-stationary solution $\mathbf{w}$ to (9) is also a nearly $O(\epsilon)$-KKT solution to the original problem (7) in the sense that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$, $\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$, $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon),\forall i=1,2,\cdots,m$.
@@
  • if $\{g_{i}\}_{i=0}^{m}$ are weakly convex, then with $\tilde{\mathcal{O}}\left(\frac{m}{B_{1}B_{2}\delta^{4}\epsilon^{5}}\right)$ iterations ALEXR2 find a nearly $\epsilon$-KKT solution $\mathbf{w}$ for (7) satisfying that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\mathbb{E}\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\mathbb{E}\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$ and it holds with probability [^1] If $|g_{i}(\mathbf{w})|<+\infty$ for any $i$ and $\mathbf{w}$, this high probability result can be replaced by $\mathbb{E}[\max\limits_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})]\leq O(\epsilon)$ and $\mathbb{E}[|g_{i}(\hat{\mathbf{w}})\nu_{i}|]\leq O(\epsilon),\forall i=1,2,\cdots,m$. $1-O(\epsilon)$ that $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon),\forall i=1,2,\cdots,m$.
@@
Notably, our method not only improve the rate but also enjoy an improved dependence on $\delta$, compared to $O(\epsilon^{-6}\delta^{-6})$ in prior work [8].
6 Numerical Experiments
We conduct experiments to verify the effectiveness of the proposed algorithms. For non-smooth non-convex FCCO, we consider GDRO with CVaR divergence [5]. For non-convex constrained optimization, we follow [8] and consider two tasks, namely AUC maximization with ROC fairness constraints and continual learning with non-forgetting constraints.
[…]
@@ 6 Numerical Experiments
We use 3 datasets: Camelyon17, Amazon [34], and CelebA [35]. The first two datasets are from WILDS Benchmark for evaluating methods tackling the distributional shift [36], where Camelyon17 has 30 groups and Amazon has 1252 groups. CelebA is a large-scale facial attribute dataset containing over 200,000 celebrity images annotated with 40 binary attributes. We select 4 binary attributes ‘Attractive’, ‘Mouth_Slightly_Open’, ‘Male’ and ‘Blonde_Hair’ and construct 16 groups, where ‘Blonde_Hair’ also serves as the label for classification. We compare SONEX with SONX and OOA, where OOA is a stochastic primal-dual algorithm  [^2] We extend the original OOA [37] to GDRO with CVaR setting as in [5]. We use pretrained Densenet121 [38], Distilbert [39] and ResNet50 [40] for Camelyon17, Amazon and CelebA, respectively. We perform Adam-type update for SONEX on Amazon dataset and momentum-type update on Camelyon17 and CelebA datasets. We run each experiment for 3 random seeds and report their average performance. The hyperparameter tuning is discussed in Appendix F.1.